Steam enters a turbine at 9 MPa and T1°C and leaves at 20 kPa with a quality equal to 0.60. Neglecting the changes in kinetic and potential energies, determine T1 if the mass flow rate through the turbine is 4 kg/s. The turbine’s power output is 7.5 MW and the heat loss from the turbine is 5 kJ/kg.
Answer #1

Turbine inlet temperature is 563.885 °Cwin Gin. Qi P = 9 MPa T, 8c Pg - 20kPa, x=0-6 Olom in in kinetic and petential energies the changes 2 Neglect za ./₂² -20 28hi = 1880 + 1665.908 = 3845-908 kJ/reg P = 9 MPa = 2000 kPa For shake hgi = 2742.9 L 35 45.908 hg, chi steam is in eeyner she